AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu May 22 06:41:10 2025
Author
*
:
Subject
*
:
07.00 JG found target off (target data stopped c. 02.30 this morning) and beam off (not re-enabled from previous CRYRNG area entry - target related?) JG re-enables beam 07.37 Run R5 stopped Run R6 started - switch to no storage mode 07.44 DSSSD bias and leakage currents - attachments 1 FEE64 Rate spectra OK - attachment 2 ADC data item stats OK - attachment 3 FEE64 temps OK - attachment 4 WR timestamps OK - attachment 5 Merger and Tape Server - attachments 6 DSSSD x-y hit patterns - attachment 8 Per JM instructions will move detectors toward beam axis by 30mm and investigate effect of e-target scrapers 08.45 Run R7 JG has setup e-target scrapers (YR09DS1HA and YR09DS1HI) both scrapers are set to move with the beam cycle. They move in with the start of the 'target' period (SC3). They move out with the start of the 'intermediate' period (SC4). The movement is rather slow (~1sec) so there is still some unblocked time during 'target' period. The positions are YR09DS1HA (outside): 10mm (SC3) and 30mm (SC4) YR09DS1HI (inside): -0.5mm (SC3) and -20mm (SC4) 11:50 Jan has been playing with the scraper settings so that we can take a run with the scrapers and without. Since he has been playing. We see the beam halo. This is when scrapers were in and beam was moved to the inside of the ring by 5mm. See attachment 10 Moving the beam to the putside of the ring by 5mm we see halo on the left now. See attachment 11 We believe this is scattering from the edge of the scrapers. See attachment 12 for the energy spectrum which corresponds to attachment 11. It is consistent with scattering. We will now take a run with and without the scrapers in and with the left detector moving in with every cycle via the pneumatic motor. 12:20 R8 started. Scrapers in. Left arm moving in synchronous with the ring cycle, potentiometer when IN = 62.71 mm. Right arm stationary and OUT, potentiometer = 19.8 mm 13:00 Run 9 started. Same condition as run 8 except no scrapers. 13:35 Run 9 stopped, R10 started -> no storage 13.37 Merger - attachment 13 per FEE64 n+n Ohmic 1.8.W spectra - 20us, 200us, 2ms, 20ms FSR - attachments 14-18 per FEE64 p+n junction 1.8.W spectra - 20ms, 2ms, 200us, 20us FSR - attachments 19-22 per FEE64 Rate spectra - attachment 23 WR timestamp OK - attachment 24 ADC, DISC, PAUSE RESUME data item stats - attachment 25-28 FEE64 temps OK - attachment 29 DSSSD bias & leakage current OK - attachment 30 14:00 R11 Pulser walkthrough See attachments 31-32 Following pulser walkthrough R12 started with no beam and no target in the ring. Cave is accessible while we switch to 7Li 20:15 System checks OK
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2025-05-22_07-44-10.png
Original size: 734x499
Attachment 2:
Screenshot_from_2025-05-22_07-44-21.png
Original size: 1920x1014
Attachment 3:
Screenshot_from_2025-05-22_07-44-30.png
Original size: 1920x1014
Attachment 4:
Screenshot_from_2025-05-22_07-44-44.png
Original size: 1920x1014
Attachment 5:
Screenshot_from_2025-05-22_07-45-17.png
Original size: 1920x1014
Attachment 6:
Screenshot_from_2025-05-22_07-45-27.png
Original size: 1920x1014
Attachment 7:
Screenshot_from_2025-05-22_07-45-36.png
Original size: 1920x1014
Attachment 8:
Screenshot_from_2025-05-22_07-46-08.png
Original size: 1920x1014
Attachment 9:
Screenshot_from_2025-05-22_07-46-18.png
Original size: 787x606
Attachment 10:
Screenshot_from_2025-05-22_11-59-16.png
Original size: 1920x1014
Attachment 11:
Screenshot_from_2025-05-22_12-09-58.png
Original size: 1920x1014
Attachment 12:
Screenshot_from_2025-05-22_12-13-06.png
Original size: 1920x1014
Attachment 13:
Screenshot_from_2025-05-22_13-47-20.png
Original size: 1920x1014
Attachment 14:
Screenshot_from_2025-05-22_13-47-07.png
Original size: 1920x1014
Attachment 15:
Screenshot_from_2025-05-22_13-46-10.png
Original size: 1920x1014
Attachment 16:
Screenshot_from_2025-05-22_13-45-40.png
Original size: 1920x1014
Attachment 17:
Screenshot_from_2025-05-22_13-45-17.png
Original size: 1920x1014
Attachment 18:
Screenshot_from_2025-05-22_13-43-26.png
Original size: 1920x1014
Attachment 19:
Screenshot_from_2025-05-22_13-42-53.png
Original size: 1920x1014
Attachment 20:
Screenshot_from_2025-05-22_13-42-23.png
Original size: 1920x1014
Attachment 21:
Screenshot_from_2025-05-22_13-41-44.png
Original size: 1920x1014
Attachment 22:
Screenshot_from_2025-05-22_13-41-14.png
Original size: 1920x1014
Attachment 23:
Screenshot_from_2025-05-22_13-39-05.png
Original size: 1920x1014
Attachment 24:
Screenshot_from_2025-05-22_13-38-52.png
Original size: 1920x1014
Attachment 25:
Screenshot_from_2025-05-22_13-38-07.png
Original size: 1920x1014
Attachment 26:
Screenshot_from_2025-05-22_13-37-56.png
Original size: 1920x1014
Attachment 27:
Screenshot_from_2025-05-22_13-37-42.png
Original size: 1920x1014
Attachment 28:
Screenshot_from_2025-05-22_13-37-25.png
Original size: 1920x1014
Attachment 29:
Screenshot_from_2025-05-22_13-37-14.png
Original size: 1920x1014
Attachment 30:
Screenshot_from_2025-05-22_13-36-40.png
Original size: 734x499
Attachment 31:
Screenshot_from_2025-05-22_14-40-46.png
Original size: 1920x1014
Attachment 32:
Screenshot_from_2025-05-22_14-40-00.png
Original size: 1920x1080
Attachment 33:
Screenshot_from_2025-05-22_15-06-01.png
Original size: 1095x846
Attachment 34:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898